OpenStudy (anonymous):

What is the margin of error for a yes/no survey of 200 people with a proportion of 65% who answered "yes"?

OpenStudy (anonymous):

A.±3.37% B.±5.06% C.±6.75% D.±8.43% E.±10.12%

OpenStudy (anonymous):

@amistre64

OpenStudy (whpalmer4):

I'm no statistics maven by any stretch of the imagination, but are you sure that's all of the information?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thats all i see

OpenStudy (whpalmer4):

Do you know (or can you find) the sample error?

OpenStudy (whpalmer4):

sorry, standard error, not sample error

OpenStudy (whpalmer4):

Ah, okay, got it. Standard error for a sample proportion is going to be \[\sqrt{\frac{p(1-p)}{n}}\]where \(p = 0.65\) and \(n = 200\)

OpenStudy (anonymous):

ummm im trying to lok it up but idk

OpenStudy (whpalmer4):

Hey, you're in this class, I've never studied this stuff, I shouldn't be finding it more quickly than you do :-)

OpenStudy (anonymous):

i actually failed algebra 2 and they gave me statistics so idk i've actaully nevver seen this stuff before

OpenStudy (anonymous):

i dont think that comes out as any of the answer choices

OpenStudy (anonymous):

im so confused...lol

OpenStudy (whpalmer4):

if we just evaluate that square root, the value is \[\sqrt{0.65(1-0.65)/200} \approx 0.0337268\]which expressed as a percentage would be 3.37%. That's one of our answers. However, 200 * that number is also one of our answers...you'd be better off with a real statistics weenie here, I think. Why don't you take away my medal so that people will look at the problem? (I think many people don't bother looking at questions where medals have already been awarded, figuring there is little chance they'll be able to get one).

OpenStudy (anonymous):

whats the other aswer that i could get?

OpenStudy (whpalmer4):

200 * 0.0337268 = 6.74. That string of digits appears in your choices, too. No doubt one (or maybe both) of them is a "trap" answer for people who don't understand the problem. I don't want to be responsible for you falling into the trap :-)

OpenStudy (anonymous):

what one do you think is the best option

OpenStudy (whpalmer4):

Find someone who understands the material better than I do.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

@amistre64

OpenStudy (whpalmer4):

I was looking at this page: http://stattrek.com/estimation/margin-of-error.aspx

OpenStudy (anonymous):

its 3.37 my teacher told me

OpenStudy (whpalmer4):

3.37%, not 3.37. Thanks for letting me know.